# Multimodal Prompt Perceiver: Empower Adaptiveness, Generalizability and Fidelity for All-in-One Image Restoration

### 05 Dec 2023 v1

@@ Abstract
Despite substantial progress, all-in-one image restoration (IR) grapples with persistent challenges in handling intricate real-world degradations. This paper introduces MPerceiver: a novel multimodal prompt learning approach that harnesses Stable Diffusion (SD) priors to enhance adaptiveness, generalizability and fidelity for all-in-one image restoration. Specifically, we develop a dual-branch module to master two types of SD prompts: textual for holistic representation and visual for multiscale detail representation. Both prompts are dynamically adjusted by degradation predictions from the CLIP image encoder, enabling adaptive responses to diverse unknown degradations. Moreover, a plug-in detail refinement module improves restoration fidelity via direct encoder-to-decoder information transformation. To assess our method, MPerceiver is trained on 9 tasks for all-in-one IR and outperforms state-of-the-art task-specific methods across most tasks. Post multitask pre-training, MPerceiver attains a generalized representation in low-level vision, exhibiting remarkable zero-shot and few-shot capabilities in unseen tasks. Extensive experiments on 16 IR tasks and 26 benchmarks underscore the superiority of MPerceiver in terms of adaptiveness, generalizability and fidelity.
@@ 1 Introduction
Image restoration (IR) aims to reconstruct a high-quality (HQ) image from its degraded low-quality (LQ) counterpart. Recent deep learning-based IR approaches excel in addressing single degradation, such as denoising [117, 118, 102], deblurring [84, 93, 78], adverse weather removal [107, 12, 34, 72, 44, 63], low-light enhancement [26, 109, 105], etc. However, these task-specific methods often fall short in real-world scenarios, such as autonomous driving and outdoor surveillance, where images may encounter unknown, dynamic degradations [66, 132]. The concept of all-in-one image restoration has recently gained significant traction, aiming to tackle multiple degradations with a unified model using a single set of pre-trained weights. Leading approaches leverage techniques such as contrastive learning [46, 11], task-specific sub-networks [73, 130], task-specific priors [107, 94], and task-agnostic priors [60] to enhance the network’s capability across various degradations. Despite their promising performance, the adaptability and generalizability of all-in-one models to real-world scenarios, characterized by intricate and diverse degradations, remain challenging.
@@
To demonstrate the superiority of MPerceiver as an all-in-one approach, it’s trained on 9 IR tasks covering both synthetic and real settings. As shown in Fig. 2, our method outperforms all compared all-in-one methods and achieves even better results than state-of-the-art task-specific methods across most tasks. Besides, MPerceiver can even handle challenging mixed degradations that may occur in real-world scenarios (Fig. 1 I). Furthermore, after multitask pre-training, MPerceiver has learned general representations in low-level vision. Comprehensive experiments show that pre-trained MPerceiver exhibits favorable zero-shot and few-shot capabilities in 6 unseen tasks (Fig. 1 II, III).
@@
  • Extensive experiments on 16 image restoration tasks (all-in-one, zero-shot, few-shot) and 26 benchmarks validate MPerceiver’s superiority in achieving high adaptiveness, robust generalizability, and superior fidelity when addressing diverse intricate degradations.
@@ 2.1 Image Restoration
Image restoration (IR) methods for known degradations [55, 112, 70, 8, 113, 100, 9, 54, 124, 116, 128] have been widely explored, while all-in-one approaches are still in the exploratory stage [51, 11, 60, 73]. TransWeather [94] designs a transformer-based network with learnable weather type queries to tackle different types of weather. AirNet [46] recovers various degraded images through a contrastive-based degraded encoder. Zhu et al. [130] propose a strategy for investigating both weather-general and weather-specific features. IDR [115] employs an ingredients-oriented paradigm to investigate the correlation among various restoration tasks. Most of the existing methods are capable of handling a limited range of degradation types and cannot cover complex real-world scenarios.
@@ 2.2 Prompt Learning for Vision Tasks
Inspired by the success of prompt learning in NLP [7, 23, 53, 62], the computer vision community has begun to explore its applicability to vision [40, 101, 37, 61] and vision-language models [127, 126, 81, 20]. CoOp [127] learns a set of task-specific textual prompts to fine-tune CLIP [79] for downstream image recognition. CoCoOP [126] refines the generalizability of CoOp by learning a lightweight neural network to generate image-conditional dynamic prompts. VPT [40] learns a set of visual prompts to finetune transformer-based vision models for downstream recognition tasks. While prevailing methods are primarily tailored for high-level vision tasks, ours is the first to explore multimodal prompt design in low-level vision.
@@ 3.3 Detail Refinement Module
While the proposed dual-branch module empowers Stable Diffusion (SD) with a robust ability to identify and eliminate degradations in images, the generated images may exhibit a tendency to lose details of small objects. This effect is attributed to the high compression rate of the Variational Autoencoder (VAE) employed by SD, as discussed in previous works [16, 133]. For instance, in Fig. 5 (b), although the degradations in the images have been largely addressed, noticeable artifacts become apparent, particularly affecting small text.
To address this concern, we introduce a Detail Refinement Module (DRM) aimed at providing supplementary information to assist the SD VAE decoder in the image reconstruction process. As depicted in Fig. 4, DRM functions as a plug-in module, enabling direct encoder-to-decoder information transformation through a skip connection. Following modulation by the visual prompt (VP) modulator, Low-Quality (LQ) features $f_{lq}$ are concatenated with the original decoder features $f_{in}$. Subsequently, a sequence of ResBlocks and SwinBlocks[55] is employed to extract auxiliary features, enhancing detail reconstruction before the residual connection. The training of the proposed DRM involves a combination of loss functions, including reconstruction (L1) loss, color loss [97], perceptual loss [42], and adversarial loss [25].
@@ 4 Experiments
Table 1: [All-in-one] Quantitative comparison with state-of-the-art task-specific methods and all-in-one methods on 9 tasks. Considering Restormer [113] and NAFNet [9] as strong general IR baselines, we train both a Restormer model and a NAFNet model under the all-in-one setting. We denote these models with symbol $(\cdot)_{A}$ for reference. Best and second best performance are in red and blue colors, respectively.
Type Method Deraining (Rain1400) Method Dehazing (Dense-Haze) Method Desnowing (Snow100K-L)
@@
Task Specific Uformer [100] 32.84 / 0.931 23.31 / 0.061 AECRNet [104] 15.80 / 0.466 277.6 / 0.601 DesnowNet [63] 27.17/ 0.898 - / -
Restormer [113] 33.68 / 0.939 20.33 / 0.050 SGID [6] 12.49 / 0.517 379.1 / 0.699 DDMSNet [119] 28.85 / 0.877  3.24 / 0.096
DRSformer [12] 33.66 / 0.939 20.06 / 0.050 DeHamer [27] 16.62 / 0.560 309.8 / 0.658 DRT [56] 29.56 / 0.892  8.15 / 0.135
@@
All in One AirNet [46] 32.36 / 0.928 22.38 / 0.058 AirNet [46] 15.32 / 0.562 248.1 / 0.587 AirNet [46] 30.14 / 0.907  3.92 / 0.105
TransWeather [94] 32.27 / 0.927 23.57 / 0.056 TransWeather [94] 15.58 / 0.569 259.8 / 0.580 TransWeather [94] 29.92 / 0.901  3.70 / 0.109
WGWS-Net [130] 32.70 / 0.930 21.10 / 0.052 WGWS-Net [130] 15.50 / 0.535 253.8 / 0.592 WGWS-Net [130] 30.27 / 0.908  2.96 / 0.099
@@
MPerceiver (Ours) 33.69 / 0.940 17.36 / 0.047 MPerceiver (Ours) 20.19 / 0.655 150.6 / 0.394 MPerceiver (Ours) 31.11 / 0.918  2.14 / 0.085
Type Method Raindrop Removal (RainDrop) Method Low-light Enhance. (LOL-v2) Method Motion Deblur (GoPro)
@@
Quan et al. [77] 31.37 / 0.918 30.56 / 0.065 LLFlow [98] 25.92 / 0.860 73.38 / 0.173 Restormer [113] 32.92 / 0.961 10.63 / 0.086
IDT [107] 31.87 / 0.931 25.54 / 0.059 SMG [109] 24.62 / 0.867 46.88 / 0.093 Stripformer [93] 33.08 / 0.962  9.03 / 0.079
@@
All in One AirNet [46] 31.32 / 0.925 33.34 / 0.073 AirNet [46] 23.47 / 0.832 45.92 / 0.123 AirNet [46] 28.31 / 0.910 15.31 / 0.122
TransWeather [94] 31.19 / 0.923 31.07 / 0.067 TransWeather [94] 27.28 / 0.865 47.20 / 0.118 TransWeather [94] 28.27 / 0.907 18.48 / 0.157
WGWS-Net [130] 31.97 / 0.930 32.33 / 0.055 WGWS-Net [130] 24.21 / 0.858 46.82 / 0.104 WGWS-Net [130] 29.07 / 0.917 14.48 / 0.127
@@
MPerceiver (Ours) 33.62 / 0.930 19.37 / 0.044 MPerceiver (Ours) 28.06 / 0.867 32.88 / 0.072 MPerceiver (Ours) 32.98 / 0.961 10.51 / 0.087
Type Method Defocus Deblur (DPDD) Method Gaussian Denoising (Urban100) Method Real Denoising (SIDD)
@@
Task Specific DRBNet [84] 25.73 / 0.791 49.04 / 0.183 SwinIR [55] 32.56 / 0.926 13.08 / 0.058 MPRNet [112] 39.71 / 0.958 49.54 / 0.200
Restormer [113] 25.98 / 0.811 44.55 / 0.178 Uformer [100] 32.24 / 0.925 13.27 / 0.068 Uformer [100] 39.89 / 0.960 47.18 / 0.198
NRKNet [78] 26.11 / 0.810 55.23 / 0.210 Restormer [113] 32.77 / 0.929 12.11 / 0.056 Restormer [113] 40.02 / 0.960 47.28 / 0.195
FocalNet [15] 26.18 / 0.808 48.82 / 0.210 ART [116] 32.81 / 0.929 12.11 / 0.056 ART [116] 39.99 / 0.960 42.38 / 0.189
All in One AirNet [46] 25.37 / 0.770 58.82 / 0.193 AirNet [46] 31.24 / 0.906 13.10 / 0.066 AirNet [46] 38.32 / 0.945 51.20 / 0.134
TransWeather [94] 25.29 / 0.765 61.66 / 0.213 TransWeather [94] 30.98 / 0.903 13.08 / 0.054 TransWeather [94] 37.65 / 0.942 50.40 / 0.141
WGWS-Net [130] 25.31 / 0.768 60.55 / 0.187 WGWS-Net [130] 31.04 / 0.905 14.08 / 0.060 WGWS-Net [130] 37.92 / 0.941 50.92 / 0.125
@@
MPerceiver (Ours) 26.06 / 0.805 46.07 / 0.190 MPerceiver (Ours) 32.88 / 0.929 12.15 / 0.056 MPerceiver (Ours) 40.05 / 0.960 41.46 / 0.190
@@
Table 3: [All-in-one] Quantitative comparison on real-world datasets of deraining, dehazing, desnowing and motion deblurring. Methods are directly applied to the LHP [29] and RealBlur-J [82] sets to evaluate generalization to real-world images. MUSS [35] is a semi-supervised deraining model trained with additional real rainy data, denoted with * for reference.
Type Deraining (LHP [29]) Dehazing (NH-HAZE [5]) Desnowing (RealSnow [130]) Motion Deblur (RealBlur-J [82]))
Type Method    PSNR $\uparrow$    SSIM $\uparrow$ Method    PSNR $\uparrow$    SSIM $\uparrow$ Method    PSNR $\uparrow$    SSIM $\uparrow$ Method    PSNR $\uparrow$    SSIM $\uparrow$
Task Specific MUSS* [35] 30.02 0.886 AECRNet [104] 19.88 0.625 MIRNetv2 [114] 31.39 0.916 MPRNet [112] 28.70 0.873
Restormer [113] 29.72 0.889 SGID [6] 16.23 0.606 ART [116] 31.05 0.913 Restormer [113] 28.96 0.879
DRSformer [12] 30.04 0.895 DeHamer [27] 20.66 0.684 Restormer [113] 31.38 0.923 Stripformer [93] 28.82 0.876
@@
All in One AirNet [46] 31.73 0.889 AirNet [46] 17.63 0.615 AirNet [46] 31.02 0.923 AirNet [46] 27.91 0.834
TransWeather [94] 29.87 0.867 TransWeather [94] 19.19 0.644 TransWeather [94] 31.13 0.922 TransWeather [94] 28.03 0.837
WGWS-Net [130] 30.77 0.885 WGWS-Net [130] 18.39 0.639 WGWS-Net [130] 31.37 0.919 WGWS-Net [130] 28.10 0.838
MPerceiver (Ours) 32.07 0.889 MPerceiver (Ours) 21.97 0.646 MPerceiver (Ours) 31.45 0.924 MPerceiver (Ours) 29.13 0.881

### 20 Mar 2024 v2

@@ Abstract
Despite substantial progress, all-in-one image restoration (IR) grapples with persistent challenges in handling intricate real-world degradations. This paper introduces MPerceiver: a novel multimodal prompt learning approach that harnesses Stable Diffusion (SD) priors to enhance adaptiveness, generalizability and fidelity for all-in-one image restoration. Specifically, we develop a dual-branch module to master two types of SD prompts: textual for holistic representation and visual for multiscale detail representation. Both prompts are dynamically adjusted by degradation predictions from the CLIP image encoder, enabling adaptive responses to diverse unknown degradations. Moreover, a plug-in detail refinement module improves restoration fidelity via direct encoder-to-decoder information transformation. To assess our method, MPerceiver is trained on 9 tasks for all-in-one IR and outperforms state-of-the-art task-specific methods across most tasks. Post multitask pre-training, MPerceiver attains a generalized representation in low-level vision, exhibiting remarkable zero-shot and few-shot capabilities in unseen tasks. Extensive experiments on 16 IR tasks underscore the superiority of MPerceiver in terms of adaptiveness, generalizability and fidelity.
@@ 1 Introduction
Image restoration (IR) aims to reconstruct a high-quality (HQ) image from its degraded low-quality (LQ) counterpart. Recent deep learning-based IR approaches excel in addressing single degradation, such as denoising [127, 128, 111], deblurring [94, 103, 88], adverse weather removal [116, 14, 37, 81, 47, 68, 36], low-light enhancement [28, 118, 114], etc. However, these task-specific methods often fall short in real-world scenarios, such as autonomous driving and outdoor surveillance, where images may encounter unknown, dynamic degradations [73, 143]. The concept of all-in-one image restoration has recently gained significant traction, aiming to tackle multiple degradations with a unified model using a single set of pre-trained weights. Leading approaches leverage techniques such as contrastive learning [49, 13], task-specific sub-networks [82, 141], task-specific priors [116, 104], and task-agnostic priors [63] to enhance the network’s capability across various degradations. Despite their promising performance, the adaptability and generalizability of all-in-one models to real-world scenarios, characterized by intricate and diverse degradations, remain challenging.
@@
To demonstrate the superiority of MPerceiver as an all-in-one approach, it’s trained on 9 IR tasks covering both synthetic and real settings. As shown in Fig. 2, our method outperforms all compared all-in-one methods and achieves even better results than state-of-the-art task-specific methods in many tasks. Besides, MPerceiver can even handle challenging mixed degradations that may occur in real-world scenarios (Fig. 1 I). Furthermore, after multitask pre-training, MPerceiver has learned general representations in low-level vision. Comprehensive experiments show that pre-trained MPerceiver exhibits favorable zero-shot and few-shot capabilities in 6 unseen tasks (Fig. 1 II, III).
@@
  • Extensive experiments on 16 IR tasks (all-in-one, zero-shot, few-shot) validate MPerceiver’s superiority in achieving high adaptiveness, robust generalizability, and superior fidelity when addressing intricate degradations.
@@ 2.1 Image Restoration
Image restoration (IR) methods for known degradations [58, 121, 79, 10, 122, 109, 11, 57, 134, 126, 138, 139, 4] have been widely explored, while all-in-one approaches are still in the exploratory stage [54, 13, 63, 82]. TransWeather [104] designs a transformer-based network with learnable weather type queries to tackle different types of weather. AirNet [49] recovers various degraded images through a contrastive-based degraded encoder. Zhu et al. [141] propose a strategy for investigating both weather-general and weather-specific features. IDR [124] employs an ingredients-oriented paradigm to investigate the correlation among various restoration tasks. Most of the existing methods are capable of handling a limited range of degradation types and cannot cover complex real-world scenarios.
@@ 2.2 Prompt Learning for Vision Tasks
Inspired by the success of prompt learning in NLP [8, 25, 56, 65], the computer vision community has begun to explore its applicability to vision [43, 110, 40, 64] and vision-language models [137, 136, 91, 22]. CoOp [137] learns a set of task-specific textual prompts to fine-tune CLIP [89] for downstream image recognition. CoCoOP [136] refines the generalizability of CoOp by learning a lightweight neural network to generate image-conditional dynamic prompts. VPT [43] learns a set of visual prompts to finetune transformer-based vision models for downstream recognition tasks. Compared to concurrent works that introduce prompt learning into IR [83, 71, 70, 66], ours is the first to explore multimodal prompt design in low-level vision.
@@ 3.3 Detail Refinement Module
While the proposed dual-branch module empowers Stable Diffusion (SD) with a robust ability to identify and eliminate degradations in images, the generated images may exhibit a tendency to lose details of small objects. This effect is attributed to the high compression rate of the autoencoder employed by SD, as discussed in previous works [18, 144]. For instance, in Fig. 5 (b), although the degradations in the images have been largely addressed, noticeable artifacts become apparent, particularly affecting small text.
To address this concern, we introduce a Detail Refinement Module (DRM) aimed at providing supplementary information to assist the SD VAE decoder in the image reconstruction process. As depicted in Fig. 4, DRM functions as a plug-in module, enabling direct encoder-to-decoder information transformation through a skip connection. Following modulation by the visual prompt (VP) modulator, Low-Quality (LQ) features $f_{lq}$ are concatenated with the original decoder features $f_{in}$. Subsequently, a sequence of ResBlocks [11] and SwinBlocks [58] is employed to extract auxiliary features, enhancing detail reconstruction before the residual connection. The training of DRM involves a combination of reconstruction (L1) loss, color loss [107], perceptual loss [45], and adversarial loss [27].
@@ 4 Experiments
Table 1: [All-in-one] Quantitative comparison with state-of-the-art task-specific methods and all-in-one methods on 9 tasks. General IR models trained under the all-in-one setting are marked with symbol $(\cdot)_{A}$. Best and second best performance are in red and blue colors, respectively. When using the self-ensemble strategy, the model is marked with “+”.
Type Method Deraining (Rain1400) Method Dehazing (Average) Method Desnowing (Snow100K-L)
@@
Task Specific Uformer [109] 32.84 / 0.931 23.31 / 0.061 AECRNet [113] 17.84 / 0.546 225.8 / 0.526 DesnowNet [68] 27.17/ 0.898 - / -
Restormer [122] 33.68 / 0.939 20.33 / 0.050 SGID [7] 14.36 / 0.562 342.9 / 0.580 DDMSNet [129] 28.85 / 0.877  3.24 / 0.096
DRSformer [14] 33.66 / 0.939 20.06 / 0.050 DeHamer [29] 18.64 / 0.622 241.6 / 0.488 DRT [59] 29.56 / 0.892  8.15 / 0.135
@@
All in One AirNet [49] 32.36 / 0.928 22.38 / 0.058 AirNet [49] 16.48 / 0.589 219.9 / 0.479 AirNet [49] 30.14 / 0.907  3.92 / 0.105
PromptIR [83] 33.26 / 0.935 22.59 / 0.058 PromptIR [83] 16.97 / 0.595 231.9 / 0.471 PromptIR [83] 30.91 / 0.913  3.79 / 0.100
DA-CLIP [70] 29.67 / 0.851 35.01 / 0.116 DA-CLIP [70] 15.01 / 0.544 224.6 / 0.468 DA-CLIP [70] 28.31 / 0.862  3.11 / 0.098
@@
MPerceiver(Ours) 33.40 / 0.937 17.82 / 0.049 MPerceiver(Ours) 20.95 / 0.644 196.8 / 0.437 MPerceiver(Ours) 31.02 / 0.916  2.31 / 0.087
MPerceiver+(Ours) 33.69 / 0.940 17.36 / 0.047 MPerceiver+(Ours) 21.08 / 0.651 190.1 / 0.422 MPerceiver+(Ours) 31.11 / 0.918  2.14 / 0.085
Type Method Raindrop Removal (RainDrop) Method Low-light Enhance. (LOL-v2-Real) Method Motion Deblur (GoPro)
@@
Quan et al. [87] 31.37 / 0.918 30.56 / 0.065 SNR-SKF [114] 21.93 / 0.842 73.70 / 0.160 Restormer [122] 32.92 / 0.961 10.63 / 0.086
IDT [116] 31.87 / 0.931 25.54 / 0.059 RQ-LLIE [67] 22.37 / 0.854 56.92 / 0.143 Stripformer [103] 33.08 / 0.962  9.03 / 0.079
@@
All in One AirNet [49] 31.32 / 0.925 33.34 / 0.073 AirNet [49] 19.69 / 0.821 55.43 / 0.151 AirNet [49] 28.31 / 0.910 15.31 / 0.122
PromptIR [83] 32.03 / 0.938 35.75 / 0.073 PromptIR [83] 21.23 / 0.860 53.92 / 0.145 PromptIR [83] 31.02 / 0.938 17.54 / 0.131
DA-CLIP [70] 30.44 / 0.880 29.38 / 0.078 DA-CLIP [70] 21.76 / 0.762 48.23 / 0.134 DA-CLIP [70] 27.12 / 0.823 16.81 / 0.136
@@
MPerceiver(Ours) 33.21 / 0.929 21.27 / 0.051 MPerceiver(Ours) 22.16 / 0.848 45.90 / 0.130 MPerceiver(Ours) 32.49 / 0.959 10.69 / 0.089
MPerceiver+(Ours) 33.62 / 0.930 19.37 / 0.044 MPerceiver+(Ours) 22.49 / 0.854 45.29 / 0.129 MPerceiver+(Ours) 32.98 / 0.961 10.51 / 0.087
Type Method Defocus Deblur (DPDD) Method Gaussian Denoising (Average) Method Real Denoising (SIDD)
@@
Task Specific DRBNet [94] 25.73 / 0.791 49.04 / 0.183 SwinIR [58] 29.60 / 0.842 58.99 / 0.146 MPRNet [121] 39.71 / 0.958 49.54 / 0.200
Restormer [122] 25.98 / 0.811 44.55 / 0.178 Restormer [122] 29.73 / 0.845 57.56 / 0.148 Uformer [109] 39.89 / 0.960 47.18 / 0.198
NRKNet [88] 26.11 / 0.810 55.23 / 0.210 ART [126] 29.79 / 0.845 58.50 / 0.141 Restormer [122] 40.02 / 0.960 47.28 / 0.195
FocalNet [17] 26.18 / 0.808 48.82 / 0.210 Xformer [125] 29.83 / 0.847 55.22 / 0.144 ART [126] 39.99 / 0.960 42.38 / 0.189
All in One AirNet [49] 25.37 / 0.770 58.82 / 0.193 AirNet [49] 28.37 / 0.801 69.36 / 0.181 AirNet [49] 38.32 / 0.945 51.20 / 0.134
PromptIR [83] 25.66 / 0.791 52.64 / 0.197 PromptIR [83] 28.82 / 0.816 63.76 / 0.170 PromptIR [83] 39.52 / 0.954 50.52 / 0.198
DA-CLIP [70] 24.91 / 0.749 57.43 / 0.201 DA-CLIP [70] 25.13 / 0.692 59.82 / 0.235 DA-CLIP [70] 34.04 / 0.824 34.56 / 0.186
@@
MPerceiver(Ours) 25.88 / 0.803 48.22 / 0.190 MPerceiver(Ours) 29.57 / 0.838 61.44 / 0.158 MPerceiver(Ours) 39.96 / 0.959 41.11 / 0.191
MPerceiver+(Ours) 26.06 / 0.805 46.07 / 0.190 MPerceiver+(Ours) 29.61 / 0.839 60.91 / 0.156 MPerceiver+(Ours) 40.05 / 0.960 41.46 / 0.190
@@
Table 3: [All-in-one] Quantitative comparison on real-world datasets of deraining, desnowing and motion deblurring. SSID [38] has no GT images. Methods are directly applied to the LHP [31] and RealBlur-J [92] sets to evaluate generalization to real-world images. MUSS [38] is a semi-supervised deraining model trained with additional real rainy data, denoted with * for reference.
Type Deraining (LHP [31]) Deraining (SSID [38]) Desnowing (RealSnow [141]) Motion Deblur (RealBlur-J [92]))
Type Method    PSNR $\uparrow$    SSIM $\uparrow$ Method    NIQE $\downarrow$  BRISQUE $\downarrow$ Method    PSNR $\uparrow$    SSIM $\uparrow$ Method    PSNR $\uparrow$    SSIM $\uparrow$
Task Specific MUSS* [38] 30.02 0.886 MUSS* [38] 3.43 28.97 MIRNetv2 [123] 31.39 0.916 MPRNet [121] 28.70 0.873
Restormer [122] 29.72 0.889 Restormer [122] 4.12 33.29 ART [126] 31.05 0.913 Restormer [122] 28.96 0.879
DRSformer [14] 30.04 0.895 DRSformer [14] 4.19 35.52 Restormer [122] 31.38 0.923 Stripformer [103] 28.82 0.876
@@
All in One AirNet [49] 31.73 0.889 AirNet [49] 3.69 30.91 AirNet [49] 31.02 0.923 AirNet [49] 27.91 0.834
TransWeather [104] 29.87 0.867 TransWeather [104] 3.96 30.94 TransWeather [104] 31.13 0.922 TransWeather [104] 28.03 0.837
WGWS-Net [141] 30.77 0.885 WGWS-Net [141] 3.71 30.79 WGWS-Net [141] 31.37 0.919 WGWS-Net [141] 28.10 0.838
MPerceiver (Ours) 32.07 0.889 MPerceiver (Ours) 3.60 30.77 MPerceiver (Ours) 31.45 0.924 MPerceiver (Ours) 29.13 0.881